OpenStudy (realitywillslapyou):

Can someone please help me with this? I really need help. I'm not good at this type of thing. Consider the line that passes through the point (3, -2) and has a slope of 2. Part 1: Write the equation of this line using point-slope form. Part 2: Using your equation from part 1, rewrite this equation in slope-intercept form. Make sure to show all of your work. Part 3: Using your equation from part 2, rewrite this equation in standard form. Make sure to show all of your work. Thank you so much!!

OpenStudy (anonymous):

start with \[y-y_1=m(x-x_1)\] with \[m=2, x_1=3,y_1=-2\]

OpenStudy (anonymous):

you get \[y-(-2)=2(x-3)\] or \[y+2=2(x-3)\] for the second part, use the distributive property to remove the parentheses, then subtract 2 from both sides to solve for \(y\)

OpenStudy (realitywillslapyou):

Okay, thank you!